Assume now that you do not think incomes will change, but that you expect  your competitor will decrease his price by 4%. Assuming that your company  does not change the price of its aircrafts, how many would you expect your  company will sell this year?


1
Expert's answer
2021-12-05T18:53:13-0500

B.

"E(AB)d=\frac{\Delta QA}{\Delta QB}"

"1.6=\frac{\Delta QA}{-4}"

"\Delta QA=-6.4"



"-6.4=\frac{Q2-Q2}{Q2+Q1}=\frac{Q2-40000}{Q2+40000}"

-7.4Q2=-296000

Q2=40000